# Beyond Single-Granularity Prompts: A Multi-Scale Chain-of-Thought Prompt Learning for Graph

### v1

@@ Abstract.
The “pre-train, prompt” paradigm, designed to bridge the gap between pre-training tasks and downstream objectives, has been extended from the NLP domain to the graph domain and has achieved remarkable progress. Current mainstream graph prompt-tuning methods modify input or output features using learnable prompt vectors. However, existing approaches are confined to single-granularity (e.g., node-level or subgraph-level) during prompt generation, overlooking the inherently multi-scale structural information in graph data, which limits the diversity of prompt semantics. To address this issue, we pioneer the integration of multi-scale information into graph prompt and propose a Multi-Scale Graph Chain-of-Thought (MSGCOT) prompting framework. Specifically, we design a lightweight, low-rank coarsening network to efficiently capture multi-scale structural features as hierarchical basis vectors for prompt generation. Subsequently, mimicking human cognition from coarse-to-fine granularity, we dynamically integrate multi-scale information at each reasoning step, forming a progressive coarse-to-fine prompt chain. Extensive experiments on eight benchmark datasets demonstrate that MSGCOT outperforms the state-of-the-art single-granularity graph prompt-tuning method, particularly in few-shot scenarios, showcasing superior performance.

### v2

# Higher-order interactions of multi-layer prompt

@@ Abstract.
The ”pre-train, prompt” paradigm has successfully evolved in representation learning. While current prompt-tuning methods often introduce learnable prompts, they predominantly treat prompts as isolated, independent components across different network layers. This overlooks the complex and synergistic higher-order interactions that exist between prompts at various hierarchical depths, consequently limiting the expressive power and semantic richness of the prompted model. To address this fundamental gap, we propose a novel framework that explicitly models the Higher-order Interactions of Multi-layer Prompt. Our approach conceptualizes prompts from different layers not as separate entities, but as a cohesive system where their inter-relationships are critical. We design an innovative interaction module that captures these sophisticated, non-linear correlations among multi-layer prompts, effectively modeling their cooperative effects. This allows the model to dynamically aggregate and refine prompt information across the network’s depth, leading to a more integrated and powerful prompting strategy. Extensive experiments on eight benchmark datasets demonstrate that our method, by leveraging these higher-order interactions, consistently surpasses state-of-the-art prompt-tuning baselines. The performance advantage is particularly pronounced in few-shot scenarios, validating that capturing the intricate interplay between multi-layer prompts is key to unlocking more robust and generalizable representation learning.